OpenStudy (anonymous):

a square plot of land has an area of 100m what is the perimeter of the land

OpenStudy (jdoe0001):

how many equal sides does a square have?

OpenStudy (anonymous):

4

OpenStudy (jdoe0001):

how many \(\bf \text{equal sides}\) does a square have?

OpenStudy (jdoe0001):

and yes, it has 4 total

OpenStudy (anonymous):

so then 400????

OpenStudy (anonymous):

?

OpenStudy (jdoe0001):

well.... it's a SQUARE.... so... how many sides that are EQUAL to each other, does a SQUARE have? 2 sides? maybe?

OpenStudy (phi):

You are given the *area* of a square... that means you can find the length of one side. if you know the length of one side, you can find the perimeter.

OpenStudy (phi):

Do you know the formula for the area of a square ?

OpenStudy (anonymous):

Yeah a= lw

OpenStudy (phi):

true, but l and w are the same for a square.

OpenStudy (phi):

how about Area= L*L (L is the length of any side) 100 = L^2 can you find L ?

OpenStudy (phi):

Do you know about square roots ?

OpenStudy (anonymous):

it doesnt give me a little picture it just is a word problem.

OpenStudy (phi):

We don't need a picture. L^2 = 100 can you find L using square root of both sides ?

OpenStudy (jdoe0001):

 \(\bf Area = x^2\qquad Area = 100m\\ \quad \\ 100=x^2\implies x=?\)

OpenStudy (arthur_ser):

\[area = 100m ^{2}\] Thus the side(α) equals:\[\alpha=\sqrt{100}m=10m\] and the perimeter equals: \[4\alpha=4*10=40m\]